Fermi National Laboratory
 

Operation Summary for June 3, 2011 - June 6, 2011

 

The Integrated Luminosity for the period from 5/30/11 to 6/6/11 was 63.63 Pb-1 (inverse picobarns).

Friday June 3

The day shift began with Operations monitoring the Tevatron (TEV) store 8791 (160.44E30) and the Recycler stash (100.1E10), with the Antiproton Source (Pbar) stacking (5.0E10), with the Main Injector (MI), Linac, and Booster operating normally, with NuMI off due to target replacement, with the MiniBooNE, Meson FTBF T-978, and NorthStar experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams.

During the day shift, at 9:15 AM, operators reported that a kicker (I:KPS5S) tripped off and caused the loss of a transfer of antiprotons from the Pbar Accumulator to the Recycler; kicker personnel worked on this problem though out the shift, which caused 2.5 hours of down time. MTA took beam from 10:15 AM to 1:30 PM. A Linac RF station (LRF4) caused a short down time.

During the evening shift, Operations halted the MiniBooNE beam at 4:56 PM for about 20 minutes while FESS personnel cleaned the MI-12 air conditioning (AC) units; Operations had been running MiniBooNE at a reduced beam rate due to issues with its air monitoring system. At 7:33 PM, the Booster low-level RF room AC unit tripped off; FESS personnel investigated and returned the unit to service; operators reestablished Booster beam at a lower intensity due to higher temperatures. Operations disabled beam to T-978 at 8 PM. Operators reported short downtimes due to problems with LRF5, BRF1, BRF8, and BRF11.

Saturday June 4

The midnight shift began with Operations monitoring store 8791 (47.9E30) and the stash (362.2E10), with Pbar stacking (10.9E10), and with MiniBooNE and NorthStar taking beam.

During the midnight shift, Operations terminated store 8791 at 1:12 AM, put the TEV into shot setup, and established store 8792 at 2:52 AM with an average initial luminosity (AIL) of 381.27E30. Operators reported that MiniBooNE beam intensities varied due to issues with their air monitoring system. Operations reported that AC problems in the Booster low-level RF room required operators to lower beam intensities to all users.

During the day shift, Operations established beam to MTA. Operators reported that high outside air temperatures caused various AC unit problems during the shift. Operators reported that an AP150 vacuum burst caused a 10-minute loss of beam to T-978.

During the evening shift, Operations disabled beam to T-978 and to MTA at 6 PM. Operations terminated store 8792 at 8:08 PM, put the TEV into shot setup, and established store 8793 at 9:32 PM with an AIL of 385.1E30.

Sunday June 5

The midnight shift began with Operations monitoring store 8793 and the stash, with Pbar stacking, and with MiniBooNE and NorthStar taking beam.

During the midnight shift, operators reported that MiniBooNE continued to run at reduced intensities due to temperatures and air-flow problems. Operators reported that they bypassed the TEV sector D3 air compressor air drying unit due to an air leak. Operations established beam to T-978 at 4:40 AM. Operations terminated beam to the Booster NorthStar experiment at 8 AM.

During the day shift, Operations terminated store 8793 at 12:23 PM, put the TEV into shot setup, and established store 8794 at 1:38 PM with an AIL of 375.1E30. MiniBooNE continued to run at reduced rates.

During the evening shift, Operations disabled beam to T-978 at 6 PM.

Monday June 6

The midnight shift began with Operations monitoring store 8794, with Pbar stacking, and with MiniBooNE taking beam at a reduce rate.

During the midnight shift, Operations terminated store 8794 at 4:41 AM, put the TEV into shot setup, and established store 8795 at 5:38 AM with an AIL of 372.66E30. Operations established beam to T-978 at 5:58 AM.

The Plans for Monday

The plans for today are to stack, stash, and store.

 

Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 06/06/2011    email Fermilab
Security, Privacy, LegalFermi National Accelerator Laboratory